# Trio-ViT: Post-Training Quantization and Acceleration for Softmax-Free Efficient Vision Transformer

### v1

@@ Abstract
Motivated by the huge success of Transformers in the field of natural language processing (NLP), Vision Transformers (ViTs) have been rapidly developed and achieved remarkable performance in various computer vision tasks. However, their huge model sizes and intensive computations hinder ViTs’ deployment on embedded devices, calling for effective model compression methods, such as quantization. Unfortunately, due to the existence of hardware-unfriendly and quantization-sensitive non-linear operations, particularly Softmax, it is non-trivial to completely quantize all operations in ViTs, yielding either significant accuracy drops or non-negligible hardware costs. In response to challenges associated with standard ViTs, we focus our attention towards the quantization and acceleration for efficient ViTs, which not only eliminate the troublesome Softmax but also integrate linear attention with low computational complexity, and propose Trio-ViT accordingly. Specifically, at the algorithm level, we develop a tailored post-training quantization engine taking the unique activation distributions of Softmax-free efficient ViTs into full consideration, aiming to boost quantization accuracy. Furthermore, at the hardware level, we build an accelerator dedicated to the specific Convolution-Transformer hybrid architecture of efficient ViTs, thereby enhancing hardware efficiency. Extensive experimental results consistently prove the effectiveness of our Trio-ViT framework. Particularly, we can gain up to and FPS under comparable accuracy over state-of-the-art ViT accelerators, as well as and DSP efficiency. Codes will be released publicly upon acceptance.
@@
Post-training quantization, hardware acceleration, Softmax-free efficient Vision Transformer.I Introduction
@@
Thanks to the powerful global information extraction capability of self-attention mechanism, Transformers have achieved great success in various natural language processing (NLP) tasks [1, 2, 3]. This success has inspired the rapid development of Vision Transformers (ViTs) [4, 5], which have gained increasing attention in the field of computer vision and shown superior performance compared to their convolution-based counterparts. However, their enormous model sizes and intensive computations challenge the deployment of ViTs on embedded/mobile devices, where both memory and computing resources are limited. For example, ViT-Large [4] contains M parameters and yields G FLOPs during inference. Thus, effective model compression techniques are highly desired to facilitate ViT’s real-world applications.
@@
Among them, model quantization [6, 7, 8] stands out as one of the most effective and widely adopted compression methods. It converts floating-point weights/activations into integers, leading to a reduction in both memory consumption and computational costs during inference. Unfortunately, due to the existence of non-linear operations, including LayerNorm (LN), GELU, and especially Softmax, which are not only hardware unfriendly but also quantization-sensitive, ViTs are difficult to be fully quantized, yielding either significant accuracy drops or notable hardware overhead [9, 10]. To solve these challenges, several efforts [9, 11, 10] have been devoted. For example, FQ-ViT [9] identifies extreme inter-channel variations in LN’s inputs and excessive non-uniform distributions in attention maps, and proposes Power-of-Two Factor (PTF) and Log-Int-Softmax (LIS) for LN and Softmax quantization, respectively. Furthermore, I-ViT [11] develops innovative lightweight dyadic arithmetic methods to approximate ViTs’ non-linear operations, thus achieving integer-only inference. Despite their effectiveness, they are dedicated to the quantization for standard ViTs while overlooking the inherent quantization and acceleration opportunities within efficient ViTs [12, 13, 14], where the vanilla Softmax-based attention with quadratic computational complexity is typically traded with more efficient Softmax-free attentions that exhibit linear computational complexity. Thus, we redirect our focus towards the exploration of effective quantization and acceleration for efficient ViTs, aiming to fully unleash their potential algorithmic benefits to win both accuracy and hardware efficiency, i.e., (i) the Softmax-free property to boost the achievable quantization performance and (ii) the linear complexity characteristic of attention to enhance inference efficiency.
@@
In addition to the algorithm level, various works have built dedicated accelerators to boost ViTs’ hardware efficiency from the hardware perspective [15, 16, 17]. For instance, Auto-ViT-Acc [16] adopts mixed quantization schemes, e.g., fixed-point and power-of-two, to quantize ViTs, and develops a dedicated accelerator to fully leverage the computational resources available on FPGAs. Moreover, ViTCoD [17] proposes pruning and polarization techniques to transform ViTs’ attention maps into denser and sparser variants, and then develops a dedicated accelerator incorporating both dense and sparse engines to simultaneously execute the above two workloads. Despite the superiority of the above accelerators in enhancing hardware efficiency, they are dedicated to standard ViTs and fall short in fully accelerating efficient ViTs [12, 13, 14], which are typically characterized by (i) Softmax-free linear attentions and (ii) Convolution-Transformer hybrid architectures. Specifically, it has been widely verified that the computational complexity reduction of linear attention yields a degradation in its local feature extraction ability, thus necessitating extra compensation components such as convolutions [12, 13]. This results in hybrid architectures for efficient ViTs that comprise both convolutions and Transformer blocks, thus calling for dedicated accelerators to unleash their potential benefits.
@@
Extensive experiments and ablation studies consistently validate the effectiveness of our Trio-ViT framework. For example, we can offer up to and FPS with comparable accuracy over state-of-the-art (SOTA) ViT accelerators. Besides, we can achieve up to and DSP efficiency. It is expected that our work can open up an exciting perspective for the quantization and acceleration of Softmax-free efficient ViTs.
The rest of this paper is organized as follows: we first introduce related works in Sec. II and preliminaries in Sec. III; Then Sec. IV and Sec. V separately illustrates Trio-ViT’s post-training quantization and dedicated accelerator; Furthermore, extensive experiments and ablation studies consistently demonstrates our Trio-ViT’s effectiveness in Sec. VI; Finally, Sec. VII summarizes this paper.
@@ II-A Model Quantization for Vision Transformers (ViTs)
Model quantization, which represents floating-point weights and activations with integers without modifying model architectures, is a generic compression solution. It can be roughly categorized into two approaches: quantization-aware training (QAT) and post-training quantization (PTQ). Specifically, QAT [18, 8, 19, 11] involves weight fine-tuning to facilitate quantization, yielding higher accuracy or lower quantization bit. In contrast, PTQ [9, 10, 20, 21], which eliminates resource-intensive fine-tuning and streamlines models’ deployment, has recently gained increasing attention. For example, [20] incorporates an innovative ranking loss to preserve the functionality of the self-attention mechanism during quantization, successfully quantizing linear operations (matrix multiplications) in ViTs. Additionally, FQ-ViT [9] further introduces Power-of-Two Factor (PTF) and Log-Int-Softmax (LIS) to quantize the hardware- and quantization-unfriendly non-linear operations (i.e., LayerNorm and Softmax) in ViTs, achieving full quantization. However, these works are developed for standard ViTs and cannot capture quantization opportunities offered by efficient ViTs [12, 13], which feature Softmax-free attention with linear computational complexity to win both quantization accuracy and hardware efficiency.
@@ II-B Efficient ViTs
ViTs [4, 5, 22, 12, 13, 23, 24] have gained growing attention recently and have been developed rapidly in the computer vision field. Among them, ViT [4] firstly applies a pure Transformer to process sequences of image patches, achieving remarkable performance. Furthermore, DeiT [5] offers a better training recipe for ViT, significantly reducing training costs. However, ViTs still incur intensive computational costs and huge memory footprints during inference to achieve superior performance, calling for efficient ViTs [22, 12, 13, 23, 24]. Particularly, EfficientViT [12], the SOTA one, replaces the vanilla self-attention of quadratic complexity with a novel lightweight multi-scale attention, achieving a global receptive field while enhancing hardware efficiency. Besides, Flatten Transformer [13] opts for an innovative focused linear attention, preserving expressiveness with low computational complexity. Despite the inherent algorithmic benefits of Softmax-free linear attentions in efficient ViTs [12, 13], including (i) the Softmax-free property to facilitate quantization and (ii) the linear complexity to boost hardware efficiency, their dedicated quantization and acceleration methods remain under-explored.
@@ II-C Transformer Accelerators
Recently, various works [25, 26, 16, 27, 17, 15] have developed dedicated accelerators to promote Transformers’ real-world deployment. Specifically, Sanger [25] dynamically prunes attention maps during inference and builds a reconfigurable accelerator with a score-stationary dataflow to accelerate such sparse patterns. As for ViT accelerators, VAQF [15] is designed to accelerate ViTs with binary weights and mixed-precision activations. Auto-ViT-Acc [16] incorporates heterogeneous computing resources available on FPGAs (i.e., DSPs and LUTs) to separately accelerate the mixed quantization scheme (i.e., fixed-point and power-of-two) for ViTs. ViTCoD [17] prunes and polarizes ViTs’ attention maps into denser and sparser ones, and constructs an accelerator to execute them on separate computing engines. While these methods can enhance the hardware efficiency for standard ViTs, they are not directly applicable for efficient ViTs [12, 13] due to the latter’s distinct model architectures, such as Softmax-free linear attention and Convolution-Transformer hybrid structures, calling for dedicated accelerators.
@@ III-A Structure of Standard ViTs
As depicted in Fig. 1, input images are initially partitioned into fixed-size patches and further enhanced with token and positional embeddings, serving as input tokens for ViTs’ Transformer blocks. Each Transformer block comprises two key components: a Multi-Head Self-Attention module (MHSA) and a Multi-Layer Perceptron (MLP), both preceded by LayerNorm (LN) and followed by residual connections. Specifically, the MHSA is the core element in Transformers for global information capture. It projects input tokens into queries , keys , and values following Eq. (1), where , , and are corresponding weights for the head. Subsequently, as formulated in Eq. (2), is multiplied by the transposed () and then subjected to Softmax normalization (where represents the feature dimension of each head) to generate the attention map. This attention map is further multiplied by to obtain the attention output for the head. Finally, the attention outputs from all heads are concatenated and projected using weights to produce the final MHSA output, i.e., in Eq. (3). As for the MLP, it comprises two linear layers separated by the GELU activation function.
Limitations. There are two limitations of standard ViTs: (i) the quadratic computational complexity of the self-attention w.r.t. token numbers [22, 12, 13] and (ii) the hardware- and quantization-unfriendly non-linear operations, i.e., LN, GELU, and especially Softmax [9, 10], hindering ViTs’ achievable hardware efficiency and quantization accuracy.
@@ III-B Structure of EfficientViT
Specifically, EfficientViT [12] mainly consists of two types of blocks: MBConvs [28] and EfficientViT modules. Each MBConv comprises two point-wise convolutions (PWConvs) divided by a depthwise convolution (DWConv), each followed by a BN and a Hswish (except the last layer). Particularly, BN can be implemented using convolutions and seamlessly folded into preceding convolutions, simplifying quantization and acceleration [30]. Additionally, each EfficientViT module includes a lightweight Multi-Scale Attention (MSA) for context information extraction and an MBConv for local information extraction. In MSA, inputs are projected to generate , which are then processed by lightweight small-kernel convolutions to generate multi-scale tokens. After applying ReLU-based global attention, the results are concatenated and projected to produce final outputs. Notably, ReLU-based global attention essentially replaces the similarity function in vanilla Softmax-based attention with , thus allowing for (i) the removal of Softmax and (ii) the utilization of the associative property of matrix multiplication to reduce computational complexity from quadratic to linear. This reformulates Eq. (2) as follows:
@@ IV Trio-ViT’s Post-Training Quantization
We adopt the most widely applied hardware-friendly quantization setting [6] by default, i.e., the symmetric layer-wise and filter-wise uniform quantization for activations and weights , respectively. Formally, as expressed in Eq. (5), / are quantized /, / are the corresponding scaling factors, means rounding to the nearest, and is the quantization bandwidth. Particularly, we follow the SOTA PTQ method BRECQ [6], which uses the diagonal Fisher Information Matrix (FIM) to sequentially reconstruct basic blocks (e.g., MBConvs and Lightweight MSAs in EfficientViT), thus enhancing cross-layer dependency while maintaining generalizability. Given the FIM as the objective function, quantization of weights and activations are optimized via Adaround [31] and Learned Step size Quantization (LSQ) [32], respectively.
@@ IV-A Observations
As we adopt the block-wise reconstruction following [6] for quantization optimization and MBConvs/lightweight MSAs are two primary blocks in EfficientViT, we begin by retaining matrix-multiplications (MatMuls) within MSAs in Eq. (4) at full precision to assess quantization impact on MBConvs.
@@ IV-A1 Observations on Quantization of MBConvs
Although it has been widely recognized that activations are more sensitive to quantization than weights [9, 21], this sensitivity is exacerbated in the context of EfficientViT. As presented in Table I, quantizing only weights in EfficientViT-B1 [12] to -bit (W8) leads to a comparable accuracy () compared to its full-precision counterpart, while quantizing both weights and activations (except MatMuls in MSAs) to the same bit (W8A8) yields a catastrophic accuracy drop of . This emphasizes the extreme quantization sensitivity of activations in EfficientViT, particularly those in MBConvs, as evident in the accuracy comparison between columns and in Table II. As previously introduced in Sec. III-B, each MBConv contains two PWConvs separated by a DWConv, we then conduct an ablation study to individually quantize input activations of these three layers in all MBConvs. As shown in Table II, the input activations of DWConvs (DW) and the second PWConvs (PW2) are the most quantization sensitive and should be primarily responsible for the accuracy drop. To comprehend this issue, we visualize their input activations in Fig. 3 and observe two challenges.
@@
Challenge # 2: Inter-Channel Asymmetries in PW2’s Inputs. As depicted in Fig. 3 (e), input activations of PW2 feature extreme inter-channel asymmetries compared to those of PW1 in Fig. 3 (a), yielding a broader value range and thus a lower quantization resolution. For instance, in PW2’s input within the MBConv from the last stage of EfficientViT-B1, the interval of the first channel is (, ), while the interval among all channels is (, ), which is larger.

### v3

@@ Abstract
Motivated by the huge success of Transformers in the field of natural language processing (NLP), Vision Transformers (ViTs) have been rapidly developed and achieved remarkable performance in various computer vision tasks. However, their huge model sizes and intensive computations hinder ViTs’ deployment on embedded devices, calling for effective model compression methods, such as quantization. Unfortunately, due to the existence of hardware-unfriendly and quantization-sensitive non-linear operations, particularly Softmax, it is non-trivial to completely quantize all operations in ViTs, yielding either significant accuracy drops or non-negligible hardware costs. In response to challenges associated with standard ViTs, we focus our attention towards the quantization and acceleration for efficient ViTs, which not only eliminate the troublesome Softmax but also integrate linear attention with low computational complexity, and propose Trio-ViT accordingly. Specifically, at the algorithm level, we develop a tailored post-training quantization engine taking the unique activation distributions of Softmax-free efficient ViTs into full consideration, aiming to boost quantization accuracy. Furthermore, at the hardware level, we build an accelerator dedicated to the specific Convolution-Transformer hybrid architecture of efficient ViTs, thereby enhancing hardware efficiency. Extensive experimental results consistently prove the effectiveness of our Trio-ViT framework. Particularly, we can gain up to , , and FPS under comparable accuracy over state-of-the-art ViT accelerators, as well as , , and DSP efficiency. Codes are available at https://github.com/shihuihong214/Trio-ViT.
@@
Post-training quantization, hardware acceleration, Transformer, Softmax-free efficient Vision Transformer.I Introduction
@@
Thanks to the powerful global information extraction capability of self-attention mechanism, Transformers have achieved great success in various natural language processing (NLP) tasks [1, 2, 3]. This success has inspired the rapid development of Vision Transformers (ViTs) [4, 5], which have gained increasing attention in the field of computer vision and shown superior performance compared to their convolution-based counterparts. However, their enormous model sizes and intensive computations challenge the deployment of ViTs on embedded/mobile devices, where both memory and computing resources are limited. For example, ViT-Large [4] contains M parameters and yields G FLOPs during inference. Thus, effective model compression techniques are highly desired to facilitate ViTs’ real-world applications.
@@
Among them, model quantization [6, 7, 8, 9] stands out as one of the most effective and widely adopted compression methods. It converts floating-point weights/activations into integers, leading to a reduction in both memory consumption and computational costs during inference. Unfortunately, due to the existence of non-linear operations, including LayerNorm (LN), GELU, and especially Softmax, which are not only hardware unfriendly but also quantization-sensitive, ViTs are difficult to be fully quantized, yielding either significant accuracy drops or notable hardware overhead [10, 11]. To solve these challenges, several efforts [10, 12, 11] have been devoted. For example, FQ-ViT [10] identifies extreme inter-channel variations in LN’s inputs and excessive non-uniform distributions in attention maps, and proposes Power-of-Two Factor (PTF) and Log-Int-Softmax (LIS) for LN and Softmax quantization, respectively. Additionally, I-ViT [12] develops innovative lightweight dyadic arithmetic methods to approximate ViTs’ non-linear operations, thus achieving integer-only inference. Despite their effectiveness, they are dedicated to the quantization for standard ViTs while overlooking the inherent quantization and acceleration opportunities within efficient ViTs [13, 14, 15], where the vanilla Softmax-based attention with quadratic computational complexity is typically traded with more efficient Softmax-free attentions that exhibit linear computational complexity. To close this gap, we redirect our focus towards the exploration of effective quantization and acceleration for efficient ViTs, aiming to fully unleash their potential algorithmic benefits to win both accuracy and hardware efficiency, i.e., (i) the Softmax-free property to boost the achievable quantization performance and (ii) the linear complexity characteristic of attentions to enhance inference efficiency.
@@
In addition to the algorithm level, various works have built dedicated accelerators to boost ViTs’ hardware efficiency from the hardware perspective [16, 17, 18]. For instance, Auto-ViT-Acc [17] adopts mixed quantization schemes, i.e., fixed-point and power-of-two, to quantize ViTs, and develops a dedicated accelerator to fully leverage the computational resources available on FPGAs. Moreover, ViTCoD [18] proposes pruning and polarization techniques to transform ViTs’ attention maps into denser and sparser variants, and then develops a dedicated accelerator incorporating both dense and sparse engines to simultaneously execute the above two workloads. Despite the superiority of the above accelerators in enhancing hardware efficiency, they are dedicated to standard ViTs and fall short in fully accelerating efficient ViTs [13, 14, 15], which are typically characterized by (i) Softmax-free linear attentions and (ii) Convolution-Transformer hybrid architectures. Specifically, it has been widely verified that the computational complexity reduction of linear attentions will yield a degradation in their local feature extraction ability, thus necessitating extra compensation components such as convolutions [13, 14]. This results in hybrid architectures for efficient ViTs that comprise both convolutions and Transformer blocks, thus calling for dedicated accelerators to unleash their potential benefits.
@@
Extensive experiments and ablation studies consistently validate the effectiveness of our Trio-ViT framework. For example, we can offer up to , , and FPS with comparable accuracy over state-of-the-art (SOTA) ViT acceleration frameworks. Besides, we can achieve up to , , and DSP efficiency. It is expected that our work can open up an exciting perspective for the quantization and acceleration of Softmax-free efficient ViTs.
The rest of this paper is organized as follows: we first introduce related works in Sec. II and preliminaries in Sec. III; Then, we illustrate Trio-ViT’s post-training quantization engine and dedicated accelerator in Sec. IV and Sec. V, respectively; Furthermore, extensive experiments and ablation studies consistently demonstrate Trio-ViT’s effectiveness in Sec. VI; Finally, Sec. VII summarizes this paper.
@@ II-A Model Quantization for Vision Transformers (ViTs)
Model quantization, which represents floating-point weights and activations with integers without modifying model architectures, is a generic compression solution. It can be roughly categorized into two approaches: quantization-aware training (QAT) and post-training quantization (PTQ). Specifically, QAT [19, 8, 20, 12] involves weight fine-tuning to facilitate quantization, yielding higher accuracy or lower quantization bit. In contrast, PTQ [10, 11, 21, 22, 9], which eliminates resource-intensive fine-tuning and streamlines models’ deployment, has recently gained increasing attention. For example, [21] incorporates an innovative ranking loss to preserve the functionality of the self-attention mechanism during quantization, successfully quantizing linear operations (matrix multiplications) in ViTs. Additionally, FQ-ViT [10] further introduces Power-of-Two Factor (PTF) and Log-Int-Softmax (LIS) to quantize the hardware- and quantization-unfriendly non-linear operations (i.e., LayerNorm and Softmax) in ViTs, achieving full quantization. However, these works are developed for standard ViTs and cannot capture quantization opportunities offered by efficient ViTs [13, 14], which feature Softmax-free attention with linear computational complexity to win both quantization accuracy and hardware efficiency.
@@ II-B Efficient ViTs
ViTs [4, 5, 23, 13, 14, 24, 25] have gained growing attention recently and have been developed rapidly in the computer vision field. Among them, ViT [4] firstly applies a pure Transformer to process sequences of image patches, achieving remarkable performance. Furthermore, DeiT [5] offers a better training recipe for ViT, significantly reducing training costs. However, standard ViTs still incur intensive computational costs and huge memory footprints during inference to achieve superior performance, calling for efficient ViTs [23, 13, 14, 24, 25]. Particularly, EfficientViT [13], the SOTA one, replaces the vanilla Softmax-based self-attention of quadratic complexity with a novel Softmax-free lightweight multi-scale attention, achieving a global receptive field while enhancing hardware efficiency. Besides, Flatten Transformer [14] opts for an innovative Softmax-based focused linear attention, preserving expressiveness with low computational complexity. Despite the inherent algorithmic benefits of Softmax-free linear attentions in efficient ViTs [13, 14], including (i) the Softmax-free property to facilitate quantization and (ii) the linear complexity to boost hardware efficiency, their dedicated quantization and acceleration methods remain under-explored.
@@ II-C Transformer Accelerators
Recently, various works [26, 27, 17, 28, 18, 16] have developed dedicated accelerators to promote Transformers’ real-world deployment. Specifically, Sanger [26] dynamically prunes attention maps during inference and builds a reconfigurable accelerator with a score-stationary dataflow to accelerate such sparse patterns. As for ViT accelerators, VAQF [16] accelerates ViTs with binary weights and mixed-precision activations. Auto-ViT-Acc [17] incorporates heterogeneous computing resources available on FPGAs (i.e., DSPs and LUTs) to separately accelerate the mixed quantization schemes (i.e., fixed-point and power-of-two) for ViTs. ViTCoD [18] prunes and polarizes ViTs’ attention maps into denser and sparser ones, and constructs an accelerator to execute them on separate computing engines. While these methods can enhance the hardware efficiency for standard ViTs, they are not directly applicable for efficient ViTs [13, 14] due to their distinct model architectures, such as Softmax-free linear attention and Convolution-Transformer hybrid structures, calling for dedicated accelerators.
@@ III-A Structure of Standard ViTs
As depicted in Fig. 1, input images are initially partitioned into fixed-size patches and further enhanced with token and positional embedding, serving as input tokens for ViTs’ Transformer blocks. Among them, each Transformer block comprises two key components: a Multi-Head Self-Attention module (MHSA) and a Multi-Layer Perceptron (MLP), both preceded by LayerNorm (LN) and followed by residual connections. Specifically, the MHSA is the core element in Transformers for global information capture. It projects input tokens into queries , keys , and values following Eq. (1), where , , and are corresponding weights for the head. Subsequently, as formulated in Eq. (2), is multiplied by the transposed (i.e., ) and then subjected to Softmax normalization (where represents the feature dimension of each head) to generate the attention map. This attention map is further multiplied by to obtain the attention output for the head. Finally, the attention outputs from all heads are concatenated and projected using weights to produce the final MHSA output, i.e., in Eq. (3). As for the MLP, it comprises two linear layers separated by the GELU activation function.
Limitations. There are two limitations of standard ViTs: (i) the quadratic computational complexity of the self-attention w.r.t. token numbers [23, 13, 14] and (ii) the hardware- and quantization-unfriendly non-linear operations, i.e., LN, GELU, and especially Softmax [10, 11], which hinder ViTs’ achievable hardware efficiency and quantization accuracy.
@@ III-B Structure of EfficientViT
Specifically, EfficientViT [13] mainly consists of two types of blocks: MBConvs [29] and EfficientViT modules. Each MBConv comprises two point-wise convolutions (PWConvs) divided by a depthwise convolution (DWConv). Each layer is followed by a BN and a Hswish (except the last layer). Particularly, BN can be implemented using convolutions and seamlessly folded into preceding convolutions, simplifying quantization and acceleration [31]. Additionally, each EfficientViT module includes a lightweight Multi-Scale Attention (MSA) for context information extraction and an MBConv for local information extraction. In MSA, inputs are projected to generate , which are then processed by lightweight small-kernel convolutions to generate multi-scale tokens. After applying ReLU-based global attention, the results are concatenated and projected to produce final outputs. Notably, ReLU-based global attention essentially replaces the similarity function in vanilla Softmax-based attention with , thus allowing for (i) the removal of Softmax and (ii) the utilization of the associative property of matrix multiplication to reduce computational complexity from quadratic to linear. This reformulates Eq. (2) as follows:
@@ IV Trio-ViT’s Post-Training Quantization
We adopt the most widely applied hardware-friendly quantization setting [6] by default, i.e., the symmetric layer-wise and filter-wise uniform quantization for activations and weights , respectively. Formally, as expressed in Eq. (5), / are quantized /, / are the corresponding scaling factors, means rounding to the nearest, and is the quantization bit-width. Particularly, we follow the SOTA PTQ method BRECQ [6], which uses the diagonal Fisher Information Matrix (FIM) to sequentially reconstruct basic blocks (e.g., MBConvs and Lightweight MSAs in EfficientViT), thus enhancing cross-layer dependency while maintaining generalizability. Given the FIM as the objective function, quantization of weights and activations are optimized via Adaround [32] and Learned Step size Quantization (LSQ) [33], respectively.
@@ IV-A Observations
As we adopt the block-wise reconstruction following [6] for quantization optimization and MBConvs/lightweight MSAs are two primary blocks in EfficientViT, we begin by retaining matrix-multiplications (MatMuls) within MSAs (i.e., the computations in Eq. (4)) at full precision to assess quantization impact on MBConvs.
@@ IV-A1 Observations on Quantization of MBConvs
Although it has been widely recognized that activations are more sensitive to quantization than weights [10, 22], this sensitivity is exacerbated in the context of EfficientViT. As presented in Table I, quantizing only weights in EfficientViT-B1 [13] to -bit (W8) leads to a comparable accuracy () compared to its full-precision counterpart, while quantizing both weights and activations (except MatMuls in MSAs) to the same bit (W8A8) yields a catastrophic accuracy drop of . This emphasizes the extreme quantization sensitivity of activations in EfficientViT, particularly those in MBConvs, as evident in the accuracy comparison between columns and in Table II. As previously introduced in Sec. III-B, each MBConv contains two PWConvs separated by a DWConv, we then conduct ablation studies to individually quantize input activations of these three layers in all MBConvs. As shown in Table II, the input activations of DWConvs (DW) and the second PWConvs (PW2) are the most quantization sensitive and should be primarily responsible for the accuracy drop. To comprehend this issue, we visualize their input activations in Fig. 3 and observe two challenges.
@@
Challenge # 2: Inter-Channel Asymmetries in PW2’s Inputs. As depicted in Fig. 3 (e), input activation of PW2 features extreme inter-channel asymmetries compared to that of PW1 in Fig. 3 (a), yielding a broader value range and thus a lower quantization resolution. For instance, in PW2’s input within the MBConv from the last stage of EfficientViT-B1, the interval of the first channel is (, ), while the interval among all channels is (, ), which is larger.